a ball thrown upwards return to the ground with the initial speed of u.the ball is at a height of 80m at two times,the time interval being 6 seconds . find u by assuming g = 10m/s

Grade:9

1 Answers

Arun
6 years ago
S = 80 metres
a = g = –10 m/s^2
using the equations
S = u*t + (1/2) gt^2
80 = u* t + (1/2) ( –10) t^2
5t^2 –ut +80  = 0
t = [u + \sqrt(u^2 –1600)  ]/10
& t = [u – \sqrt(u^2 –1600)  ]/10
[u + \sqrt(u^2 –1600)  ]/10 – [u  –\sqrt(u^2 –1600)  ]/10  = 6
hence
\sqrt(u^2 –1600) = 30
(u^2 –1600) = 900
u^2 = 2500
u = 50 m/s
hope it helps